OpenStudy (anonymous):

Prove the following inequality for any triangle of sides a,b,c. -1 < a/b + b/c + c/a - (b/a + a/c + c/b) <1

OpenStudy (anonymous):

\[-1 < \frac{ a }{ b }+\frac{b }{ c }+\frac{ c }{a }-(\frac{ b }{a }+\frac{ a }{c }+\frac{c }{ c }) <1\]

OpenStudy (anonymous):

Please help. I need the answer for tomorrow. Should I use Sine Law? Thanks.

OpenStudy (goformit100):

yes you can use

OpenStudy (anonymous):

I've tried it and I'm stuck. How should I go about it?

OpenStudy (goformit100):

OK